OpenStudy (anonymous):

The equation, y = 76(1.013)x, represents the the population of the United States, after 1900, in millions. The population in 1900 (x = 0) was..? A. 1.0123 million B. 76 million C. 76.988 million D. 77 million

OpenStudy (ghazi):

well substitute x=0 you'll get B but there must be a plus sign in your equation it should be y= 76+(1.013)x

OpenStudy (anonymous):

i think it is \[ y = 76(1.013)^x\]which means in 1900 it was 76 million since \[b^0=1\] and so \[76(1.013)^0=76\times 1=76\]

OpenStudy (ghazi):

yes it is also one of the possibilities ...